### v1

# Whisper D-SGD: Correlated Noise Across Agents for Differentially Private Decentralized Learning

Abstract
Decentralized learning enables distributed agents to train a shared machine learning model through local computation and peer-to-peer communication. Although each agent retains its dataset locally, the communication of local models can still expose private information to adversaries. To mitigate these threats, local differential privacy (LDP) injects independent noise per agent, but it suffers a larger utility gap than central differential privacy (CDP). We introduce Whisper D-SGD, a novel covariance-based approach that generates correlated privacy noise across agents, unifying several state-of-the-art methods as special cases. By leveraging network topology and mixing weights, Whisper D-SGD optimizes the noise covariance to achieve network-wide noise cancellation. Experimental results show that Whisper D-SGD cancels more noise than existing pairwise-correlation schemes, substantially narrowing the CDP-LDP gap and improving model performance under the same privacy guarantees.
I Introduction
Training machine learning models traditionally involves centralizing data on a single server, which raises scalability and privacy risks [1]. To address these issues, federated learning allows distributed agents (e.g., personal devices, organization, or computing nodes in a network) to retain their data locally while sharing only data-dependent computations (e.g., local gradients or models) with a central server [2, 3]. Decentralized learning further removes the need for a central server by allowing agents to update local models and mix them directly with neighbors according to predefined mixing weights [4, 5, 6, 7]. Although federated and decentralized learning algorithms avoid sharing raw data over the network, local models may still expose sensitive information to adversaries–i.e., external eavesdroppers–through attacks, e.g., membership inference or gradient inversion [8].
A widely adopted theoretical framework for mitigating such threats is differential privacy (DP), which involves injecting random noise into data-dependent computations, under strong, formal privacy guarantees [9]. While the privacy-utility trade-off of DP has been extensively studied in the central DP (CDP) setting–where a server injects noise on the central model [10, 11, 12]–adapting DP to decentralized settings proves substantially more challenging. Under the local DP (LDP) setting [13, 14], each agent individually adds noise to its local model before mixing, protecting its local dataset from eavesdroppers. However, LDP demands substantially more noise than CDP to attain the same privacy guarantees [15].
Recent works on differentially private decentralized learning seek to reconcile the utility of CDP with the local privacy of LDP by combining (i) a small amount of independent noise per agent, sufficient to ensure local privacy, and (ii) a large amount of correlated noise among agents, designed to cancel out during mixing [16, 17, 18]. In principle, this cancellation preserves local privacy while reducing the overall noise after mixing, such that the learned models approach CDP-level utility. Among these works, [16, 17] rely on zero-sum correlation, which is well-suited for addressing honest-but-curious internal neighbors, but ineffective against external eavesdroppers as no privacy noise persists in the final model. Meanwhile, [18] proposes pairwise-canceling correlated noise, which restricts noise correlation to pairs of neighboring agents, overlooking how local models are mixed across the network and limiting the potential for broader correlation design.
In this paper, we argue that extending beyond pairwise, neighbor-only correlations can significantly improve the privacy-utility trade-off in decentralized learning. The key contributions of this paper are as follows:
  • We present a novel, covariance-based framework for the design of correlated noise across agents, that recovers state-of-the-art approaches as special cases;
  • We generalize existing analyses on the privacy-utility tradeoff, showing that the optimal covariance matrix must account for the network topology and mixing weights. Building on this insight, we propose the Whisper D-SGD algorithm, which optimizes the noise covariance to reconcile CDP-level utility with LDP-level privacy.
Experiments with various privacy budgets and network connectivity levels show that Whisper D-SGD amplifies noise cancellation over pairwise approaches, achieving superior privacy-utility trade-offs, particularly in weakly connected networks.
II Privacy-Preserving Decentralized Learning
We consider $n$ agents, $\mathcal{V}=\{1,\dots,n\}$, each with a local dataset $\mathcal{D}_{i}$, aiming to learn the parameters $x\in\mathbb{R}^{d}$ of a shared machine learning model, by minimizing the global objective:
$F(x)\triangleq\frac{1}{n}\sum_{i=1}^{n}\left[F_{i}(x)\triangleq\frac{1}{|\mathcal{D}_{i}|}\sum_{\xi_{i}\in\mathcal{D}_{i}}\ell(x,\xi_{i})\right],$ (1)
where $\ell(x,\xi_{i})$ is the loss of parameter $x$ on sample $\xi_{i}\in\mathcal{D}_{i}$ and $F_{i}(x)$ is the local objective, known only to agent $i$. Agents communicate over a network modeled by an undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where an edge $(i,j)\in\mathcal{E}$ indicates that agents $i$ and $j$ are neighbors, i.e., they can directly communicate. For simplicity, we present the case of scalar parameters ($d=1$), with straightforward extension to vector parameters.
II-A Decentralized Stochastic Gradient Descent (D-SGD)
Problem (1) is commonly solved via decentralized optimization algorithms like Decentralized Stochastic Gradient Descent (D-SGD) [6]. Each iteration $t\in\{1,\dots,T\}$ involves two steps:
a) Local step: Each agent $i$ samples $\xi_{i}^{(t)}\sim\mathcal{D}_{i}$, computes stochastic gradient $g_{i}^{(t)}=\ell^{\prime}(x_{i}^{(t)},\xi_{i}^{(t)})$, and updates parameter:
$\textstyle x_{i}^{(t+\frac{1}{2})}=x_{i}^{(t)}-\eta_{t}g_{i}^{(t)},$ (2a)
where $\eta_{t}$ is the step-size.
b) Mixing step: Agents exchange parameters with neighbors and compute a weighted average:
${x}_{i}^{(t+1)}=\sum_{j=1}^{n}w_{ij}x_{j}^{(t+\frac{1}{2})},$ (2b)
where $w_{ij}\triangleq[\mathbf{W}]_{ij}$ are mixing weights defined by the mixing matrix $\mathbf{W}\in\mathbb{R}^{n\times n}$, with $w_{ij}=0$ whenever $(i,j)\notin\mathcal{E}$. Convergence conditions for D-SGD require $\mathbf{W}$ to be doubly stochastic ($\mathbf{W}\mathbf{1}=\mathbf{1}$ and $\mathbf{1}^{\top}\mathbf{W}=\mathbf{1}^{\top}$) and the second largest eigenvalue $\lambda_{2}(\mathbf{W}^{\top}\mathbf{W})$ strictly below one [19, Section II.B].
II-B Differentially Private D-SGD
The DP framework protects the final models $\{x_{i}^{(T)}\}_{i=1}^{n}$ against adversaries (e.g., external eavesdroppers). At each iteration $t$, agent $i$ clips its local gradient $g_{i}^{(t)}$ to norm $C$, then adds privacy noise $v_{i}^{(t)}$:
$\tilde{g}_{i}^{(t)}=\hat{g}_{i}^{(t)}+v_{i}^{(t)},\quad\textstyle\hat{g}_{i}^{(t)}=\min\left\{1,\frac{C}{|g_{i}^{(t)}|}\right\}\cdot g_{i}^{(t)}.$ (3)
The resulting update is $\tilde{x}_{i}^{(t+1)}=\sum_{j=1}^{n}w_{ij}(\tilde{x}_{j}^{(t)}-\eta_{t}\tilde{g}_{j}^{(t)})$.
By relating $v_{i}^{(t)}$ to privacy parameters $(\varepsilon,\delta)$, the DP framework guarantees that the final models $\{\tilde{x}_{i}^{(T)}\}_{i=1}^{n}$, trained on any adjacent datasets $\mathcal{D},\mathcal{D}^{\prime}$ (differing in at most one agent’s dataset) satisfy, for any measurable set $\mathcal{S}$ in the parameter space:
$\mathbb{P}\bigl{[}\tilde{x}_{i}^{(T)}(\mathcal{D})\in\mathcal{S}\bigr{]}\leq e^{\varepsilon}\mathbb{P}\bigl{[}\tilde{x}_{i}^{(T)}(\mathcal{D}^{\prime})\in\mathcal{S}\bigr{]}+\delta.$ (4)
Common designs for $v_{i}^{(t)}$ include:
  1. 1.
    LDP [13, 14]. This is the standard approach, with privacy noise $v_{i}^{(t)}\sim\mathcal{N}(0,\sigma_{\text{ldp}}^{2})$ independent across agents;
  2. 2.
    DECOR [18]. The privacy noise is designed as $v_{i}^{(t)}=u_{i}^{(t)}+\sum_{j\in\mathcal{N}_{i}}v_{ij}^{(t)}$, where $u_{i}^{(t)}\sim\mathcal{N}(0,\sigma_{\text{pair}}^{2})$ is independent local noise, and each pair of neighbors $(i,j)\in\mathcal{E}$ shares a correlated noise term $v_{ij}^{(t)}=-v_{ji}^{(t)}\sim\mathcal{N}(0,\sigma_{\text{cor}}^{2})$ that cancel during mixing. Distinct pairs of neighbors produce mutually uncorrelated components, so that $v_{ij}^{(t)}$ is independent of $v_{k\ell}^{(t)}$ whenever $\{i,j\}\cap\{k,\ell\}=\emptyset$.
III A Covariance-Based Framework for
Generating Correlated Noise Across Agents
We reinterpret the differentially private D-SGD framework from a noise-covariance perspective. Instead of drawing noise components $\{v_{i}^{(t)}\}_{i=1}^{n}$ independently or from pairwise correlations, we allow agents to sample from a multivariate Gaussian:
$\mathbf{v}^{(t)}\;=\;\bigl{[}v_{1}^{(t)},\dots,v_{n}^{(t)}\bigr{]}^{\top}\;\sim\;\mathcal{N}\!\bigl{(}\mathbf{0},\mathbf{R}\bigr{)},$ (5)
where $\mathbf{R}\in\mathbb{R}^{n\times n}$ is an arbitrary covariance matrix. By defining $\tilde{\mathbf{x}}^{(t)}\triangleq[\tilde{x}_{1}^{(t)},\dots,\tilde{x}_{n}^{(t)}]^{\top}$ and $\tilde{\mathbf{g}}^{(t)}=[\tilde{g}_{1}^{(t)},\ldots,\tilde{g}_{n}^{(t)}]^{\top}$, the perturbed model-update at iteration $t$ becomes:
$\tilde{\mathbf{x}}^{(t+1)}=\mathbf{W}\left(\tilde{\mathbf{x}}^{(t)}-\eta_{t}\tilde{\mathbf{g}}^{(t)}\right).$ (6)
This covariance-based approach offers several advantages: (i) it allows for more flexible correlation structures with noise cancellation beyond pairs of neighbors; (ii) it is directly applicable to directed graphs, unlike pairwise correlation; and (iii) it recovers state-of-the-art approaches as special cases:
  1. 1.
    LDP. With independent noise, the covariance matrix is $\mathbf{R}=\sigma_{\text{ldp}}^{2}\mathbf{I}_{n}$, where $\mathbf{I}_{n}\in\mathbb{R}^{n\times n}$ is the identity matrix;
  2. 2.
    DECOR. The independent noise term $u_{i}^{(t)}\sim\mathcal{N}(0,\sigma_{\text{pair}}^{2})$ contributes $\sigma_{\text{pair}}^{2}\mathbf{I}_{n}$ to $\mathbf{R}$. In addition, each edge $(i,j)\in\mathcal{E}$ adds a pairwise correlated component $v_{ij}^{(t)}=-v_{ji}^{(t)}$ with variance $\sigma_{\text{cor}}^{2}$. In matrix form, each edge contributes $\sigma_{\text{cor}}^{2}(\mathbf{e}_{i}-\mathbf{e}_{j})(\mathbf{e}_{i}-\mathbf{e}_{j})^{\top}$, where $\mathbf{e}_{i}$ is the $i$-th standard basis vector, and summing over all edges yields $\sigma_{\text{cor}}^{2}\mathbf{L}$, where $\mathbf{L}=\sum_{(i,j)\in\mathcal{E}}(\mathbf{e}_{i}-\mathbf{e}_{j})(\mathbf{e}_{i}-\mathbf{e}_{j})^{\top}$ is the undirected graph Laplacian. The total covariance is $\mathbf{R}=\sigma_{\text{pair}}^{2}\mathbf{I}_{n}+\sigma_{\text{cor}}^{2}\mathbf{L}$.
Remark 1.
The privacy noise generation mechanism $\mathbf{v}^{(t)}\sim\mathcal{N}(\mathbf{0},\mathbf{R})$ can be implemented in a distributed fashion. A common practice is to share $\mathbf{R}$ and a random seed $s$ among the agents. At iteration $t$, each agent independently samples $\mathbf{s}^{(t)}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{n})$ using the random seed $s$ and then locally computes $\mathbf{v}^{(t)}=\mathbf{R}^{1/2}\mathbf{s}^{(t)}$.
The covariance matrix $\mathbf{R}$ controls both the level of privacy (in terms of $(\varepsilon,\delta)$-DP guarantees) and the model-update variance (hence convergence speed or utility). Intuitively, the noise must be “sufficiently large” to ensure LDP-level privacy, yet “sufficiently small” to achieve CDP-level utility.
IV Privacy-Utility Analysis
In this section, we study how the correlated noise in our general framework affects both convergence (utility) performance and $(\varepsilon,\delta)$-DP (privacy) level. Our analysis shows the interplay between noise covariance $\mathbf{R}$ and mixing matrix $\mathbf{W}$, yielding new insights into optimal noise design strategies.
IV-A Utility Analysis
To analyze utility, we introduce a virtual sequence $\hat{\mathbf{x}}^{(t+1)}=\mathbf{W}(\tilde{\mathbf{x}}^{(t)}-\eta_{t}\hat{\mathbf{g}}^{(t)})$, which begins at the noisy iterate $\tilde{\mathbf{x}}^{(t)}$ but applies a noise-free D-SGD step. This construction isolates the error introduced by the privacy noise and simplifies previous analyses (e.g., [18, Theorem 2]).
Assumption 1.
At iteration $t$, the privacy noise $\mathbf{v}^{(t)}$ is independent of the noise-free state $\hat{\mathbf{x}}^{(t+1)}$ and prior states $\mathcal{F}^{(t)}=\{(\mathbf{v}^{(1)},\hat{\mathbf{x}}^{(2)}),\dots,(\mathbf{v}^{(t-1)},\hat{\mathbf{x}}^{(t)})\}$: $\mathbb{E}[\mathbf{v}^{(t)}|{\hat{\mathbf{x}}^{(t+1)}},\mathcal{F}^{(t)}]=\mathbf{0}.$
Theorem 1.
Under Assumption 1, the expected squared error between consecutive iterates of our covariance-based framework satisfies:
$\mathbb{E}\Bigl{[}\bigl{\|}\tilde{\mathbf{x}}^{(t+1)}-\tilde{\mathbf{x}}^{(t)}\bigr{\|}^{2}\;\Big{|}\;\mathcal{F}^{(t)}\Bigr{]}$
[…]
Proof Sketch.
First, decompose $\|\tilde{\mathbf{x}}^{(t+1)}-\tilde{\mathbf{x}}^{(t)}\|^{2}$ into the sum of $\|\hat{\mathbf{x}}^{(t+1)}-\tilde{\mathbf{x}}^{(t)}\|^{2}$ and $\|\tilde{\mathbf{x}}^{(t+1)}-\hat{\mathbf{x}}^{(t+1)}\|^{2}$. Taking conditional expectation, the cross-term vanishes by Assumption 1. Next, observe $\|\tilde{\mathbf{x}}^{(t+1)}-\hat{\mathbf{x}}^{(t+1)}\|^{2}=\eta_{t}^{2}\|\mathbf{W}\mathbf{v}^{(t)}\|^{2}$. Finally, $\mathbf{R}=\mathbb{E}[\mathbf{v}^{(t)}(\mathbf{v}^{(t)})^{\top}]$ yields the variance term $\eta_{t}^{2}\,\mathrm{Tr}(\mathbf{W}\mathbf{R}\mathbf{W}^{\top})$. ∎
Theorem 1 shows that each iteration of our covariance-based framework incurs a baseline D-SGD error (independent of $\mathbf{R}$) plus a variance term proportional to $\mathrm{Tr}(\mathbf{W}\mathbf{R}\mathbf{W}^{\top})$. [^2] Due to space limits, we defer the full characterization of global convergence guarantees after $T$ iterations to an extended version of this paper.
Our analysis captures the interplay between $\mathbf{W}$ and $\mathbf{R}$, explaining how correlated noise propagates and potentially cancels across agents. Prior analyses overlooked this mutual dependency, which is central to our covariance-based view.
Another remark from Theorem 1 is that the variance $\mathrm{Tr}(\mathbf{W}\mathbf{R}\mathbf{W}^{\top})$ is linear in $\mathbf{R}$: smaller $[\mathbf{R}]_{ii}$ components speed-up convergence. At extreme cases, one may try to minimize this variance by enforcing $\mathrm{Tr}(\mathbf{W}\mathbf{R}\mathbf{W}^{\top})=0$:
(i) If $\mathbf{W}$ is invertible, then $\mathbf{R}=\mathbf{0}$ is the only solution, implying no added noise—and hence no privacy.
(ii) If $\mathbf{W}$ is singular, there exists a nontrivial null space: $\mathbf{W}\,\mathbf{z}=\mathbf{0}$, $\mathbf{z}\neq\mathbf{0}$. One can then construct $\mathbf{R}=\mathbf{z}\,\mathbf{z}^{\top}\neq\mathbf{0}$ such that $\mathbf{W}\mathbf{R}\mathbf{W}^{\top}=\mathbf{0}$; however, $\mathbf{R}$ becomes singular. Our privacy analysis will show that invertibility of $\mathbf{R}$ is necessary for general $(\varepsilon,\delta)$-DP guarantees.
IV-B Privacy Analysis
We show that our framework ensures agent-level $(\varepsilon,\delta)$-DP for arbitrary covariance $\mathbf{R}\succ 0$. The following theorem bounds $\varepsilon$ in terms of the number of iterations $T$, clipping threshold $C$, and the inverse covariance $\mathbf{R}^{-1}$.
Theorem 2.
Our covariance-based framework satisfies ($\varepsilon,\delta$)-DP with:
$\textstyle\varepsilon\leq 2C^{2}T\max\limits_{i\in\mathcal{V}}[\mathbf{R}^{-1}]_{ii}+2C\sqrt{2T\log(\frac{1}{\delta})\max\limits_{i\in\mathcal{V}}[\mathbf{R}^{-1}]_{ii}}.$ (8)
Proof Sketch.
We adapt existing Rényi DP (RDP) arguments from [12, 18] to arbitrary $\mathbf{R}\succ 0$. At each iteration, our framework satisfies $(\alpha,\alpha\varepsilon_{\text{step}})$-RDP with $\varepsilon_{\text{step}}=2C^{2}\max_{i\in\mathcal{N}}[\mathbf{R}^{-1}]_{ii}$ [18, Thm. 6]. Since RDP composes additively over $T$ iterations [12, Prop. 1], the final iterate $\tilde{\mathbf{x}}^{(T)}$ satisfies ($\alpha,T\alpha\varepsilon_{\text{step}}$)-RDP. Converting RDP to $(\varepsilon,\delta)$-DP introduces an additional term $\ln(1/\delta)/(\alpha-1)$ [12, Prop. 3]. Optimizing the expression over $\alpha>1$ concludes the proof. ∎
Theorem 2 directly relates the diagonal elements of $\mathbf{R}^{-1}$, specifically $\max_{i\in\mathcal{V}}[\mathbf{R}^{-1}]_{ii}$, to the per-agent privacy budget $\varepsilon$. This result immediately recovers well-known LDP bounds (e.g., [12, Prop. 7]) when $\mathbf{R}=\sigma_{\text{ldp}}^{2}\mathbf{I}_{n}$: larger $[\mathbf{R}]_{ii}$ (i.e., larger variance of privacy noise) implies smaller $[\mathbf{R}^{-1}]_{ii}$, thus smaller $\varepsilon$ (that is, strengthened $(\varepsilon,\delta)$-DP guarantees).
Invertibility ($\mathbf{R}\succ 0$) is a necessary condition for ensuring agent-level privacy under correlated noise: if $\mathbf{R}$ were singular, linear combinations of noise could cancel completely among agents, directly exposing their local models.
Both the number of iterations $T$ and the clipping threshold $C$ appear as multiplicative factors in Eq. (8), increasing $\varepsilon$. Intuitively, more rounds of communication or higher gradient sensitivity (larger $C$) amplify privacy loss, requiring more noise to protect local models.
V The “Whisper D-SGD” Algorithm
V-A Privacy-Utility Trade-Off
Theorems 1 and 2 together highlight a fundamental tension between utility and privacy: smaller diagonal entries $[\mathbf{R}]_{ii}$ reduce the model-update variance $\mathrm{Tr}(\mathbf{W}\mathbf{R}\mathbf{W}^{\top})$ in Eq. (7) (expediting convergence), yet they increase $[\mathbf{R}^{-1}]_{ii}$ and thus amplify $\varepsilon$ in Eq. (8) (weaker privacy). Conversely, making $[\mathbf{R}]_{ii}$ large improves privacy but slows convergence. A careful choice of $\mathbf{R}$ is therefore crucial to optimize this trade-off. From our analysis, we draw the following guidelines:
(G1) Invertibility. A necessary condition for LDP-level privacy is $\mathbf{R}\succ 0$ (Theorem 2). For this reason, we partition $\mathbf{R}$ into an independent noise component ($\sigma_{\text{mix}}^{2}\mathbf{I}_{n}$), that ensures invertibility, and an arbitrary correlated component ($\mathbf{R}_{\text{cor}}$):
$\mathbf{R}_{\text{mix}}=\sigma_{\text{mix}}^{2}\mathbf{I}_{n}+\mathbf{R}_{\text{cor}}.$ (9)
(G2) Optimization Problem. We optimize $\mathbf{R}_{\text{mix}}$ to balance utility (Theorem 1) and $(\varepsilon,\delta)$-LDP privacy (Theorem 2):
$\underset{\mathbf{R}_{\text{mix}}\succ 0}{\text{minimize}}\quad$ $\mathrm{Tr}\bigl{(}\mathbf{W}\mathbf{R}_{\text{mix}}\mathbf{W}^{\top}\bigr{)}$ (10a)
subject to $[\mathbf{R}_{\text{mix}}^{-1}]_{ii}\leq\frac{\varepsilon^{2}}{16C^{2}T\log(1/\delta)},~{}\forall\,i\in\mathcal{V}.$ (10b)
Problem (10) can be written as a convex semidefinite program (SDP). The objective is linear in $\mathbf{R}_{\text{mix}}$, and the mapping $\mathbf{R}_{\text{mix}}\mapsto\mathbf{R}_{\text{mix}}^{-1}$ is matrix-convex on $\mathbf{R}_{\text{mix}}\succ 0$ [23, §4.6.2]. By applying Schur-complement constraints, each bound on $[\mathbf{R}_{\text{mix}}^{-1}]_{ii}$ becomes a linear matrix inequality, allowing solution by standard SDP packages (e.g., MOSEK [24]). Solving (10) yields $\sigma_{\text{mix}}^{2,\star}$ and $\mathbf{R}_{\text{cor}}^{\star}$. We then compute $\mathbf{R}_{\text{mix}}^{\star}=\sigma_{\text{mix}}^{2,\star}\mathbf{I}_{n}+\mathbf{R}_{\text{cor}}^{\star}$.
Proposition 1.
Let $\varepsilon\leq\log(1/\delta)$. Problem (10) maximizes the utility of our covariance-based framework under $(\varepsilon,\delta)$-DP.
Proof.
First, the objective in (10a) minimizes the privacy-noise variance in Theorem 1, Eq. (7). Second, applying Eq. (10b) to the privacy bound in Theorem 2, Eq. (8) shows that our framework is ($\varepsilon^{*},\delta$)-DP, where $\varepsilon^{*}\leq\frac{\varepsilon^{2}}{8\log(1/\delta)}+\frac{\varepsilon}{\sqrt{2}}\leq\varepsilon$. ∎
V-B Whisper D-SGD
Algorithm 1 Whisper D-SGD
1:Input: mixing matrix $\mathbf{W}$, $(\varepsilon,\delta)$-DP parameters, number of iterations $T$, clipping threshold $C$, random seed $s$.
2:Initialize $\!\mathbf{R}_{\text{mix}}=\sigma_{\text{mix}}^{2}\mathbf{I}_{n}+\mathbf{R}_{\text{cor}}$triangleright guideline (G1)
3:Optimize (10) to find $\mathbf{R}_{\text{mix}}^{\star}$triangleright guideline (G2)
4:Share ($\mathbf{R}_{\text{mix}}^{\star}$, $s$) among all agents
5:for $t$ in $10\dots T-$ 1 do
6:     $\hat{\mathbf{g}}^{(t)}=\mathrm{clip}\bigl{(}\mathbf{g}^{(t)};C\bigr{)}$triangleright clipped gradient
7:     $\mathbf{v}^{(t)}=\mathbf{R}_{\text{mix}}^{\star 1/2}\mathbf{s}^{(t)}$, $\mathbf{s}^{(t)}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{n})$triangleright privacy noise
8:     $\tilde{\mathbf{g}}^{(t)}=\hat{\mathbf{g}}^{(t)}+\mathbf{v}^{(t)}$triangleright perturbed gradient
9:     $\tilde{\mathbf{x}}^{(t+1)}=\mathbf{W}\left(\tilde{\mathbf{x}}^{(t)}-\eta_{t}\tilde{\mathbf{g}}^{(t)}\right)$triangleright perturbed update
10:end for
11:return $\tilde{\mathbf{x}}^{(T)}$
We now introduce Whisper D-SGD (Algorithm 1), a novel differentially private D-SGD algorithm which combines our covariance-based framework and our theoretical guidelines to achieve near-CDP utility and LDP-level privacy.
Conceptual overview
In Whisper D-SGD, each agent contributes a “whisper” of noise (with variance $\sigma_{\text{mix}}^{2,\star}$) sufficient to protect its local model, thus to ensure LDP-level privacy. By design, correlated noise (with covariance $\mathbf{R}_{\text{cor}}^{\star}$) cancels during mixing, so that the final models reap many of the performance benefits typically associated with CDP-level approaches.
Implementation details
(i) Initialize $(\sigma_{\text{mix}}^{2},\mathbf{R}_{\text{cor}})$ so that $\mathbf{R}_{\text{mix}}$ is invertible (line 2). (ii) Optimize Problem (10) via standard SDP solvers (line 3). (iii) Each agent shares the optimal $\mathbf{R}_{\text{mix}}^{\star}$ and a global random seed $s$ (line 4). (iv) Execute the standard differentially-private D-SGD algorithm (lines 510), except agents add privacy noise $\mathbf{v}^{(t)}=\mathbf{R}_{\text{mix}}^{\star 1/2}\mathbf{s}^{(t)}$, where $\mathbf{s}^{(t)}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{n})$ is locally generated from seed $s$ (line 7).
Practical considerations
Once agents share $\mathbf{R}_{\text{mix}}^{\star}$ and seed $s$, Whisper D-SGD runs fully decentralized, incurring no extra communication overhead than baseline D-SGD (lines 510). However, computing $\mathbf{R}_{\text{mix}}^{\star}$ requires either a centralized or a distributed procedure (lines 24). In the centralized approach, a single node with knowledge of $\mathbf{W}$ and parameters $(\varepsilon,\delta,T,C)$ solves (10) and broadcasts $(\mathbf{R}_{\text{mix}}^{\star},s)$ to all agents. Alternatively, a distributed setup can achieve the same outcome by running an initial consensus phase on $\mathbf{W}$ and $s$, then locally solving (10). This approach removes reliance on any single node but incurs a one-time communication overhead.
Further extensions
While Whisper D-SGD outperforms pairwise and fully independent approaches against external eavesdroppers (Sections V-CVI), it currently requires sharing $(\mathbf{R}_{\text{mix}}^{\star},s)$ among all agents, offering no defense against honest-but-curious internal participants. Future work may distribute different seeds among subgroups of agents or impose structural constraints on $\mathbf{R}_{\text{cor}}$ to relax global trust. Another direction is the joint optimization of $\mathbf{W}$ and $\mathbf{R}$–rather than fixing $\mathbf{W}$–adapting the mixing weights to further amplify noise cancellation and improve convergence speed. Our framework can also be extended to directed graphs (e.g., by defining asymmetric Laplacians) and to other D-SGD variants (e.g., the push-sum algorithm [7]). Finally, if an agent departs mid-training, $\mathbf{R}_{\text{mix}}^{\star}$ must be currently recomputed. Resilient algorithms to update $\mathbf{R}_{\text{mix}}$ on the fly remain an open challenge.
V-C Comparison with Related Works
(a) $p=0.5$.
(b) $\varepsilon=10$.
Figure 1: Comparison of variances $\sigma_{\text{mix}}^{2}$, $\sigma_{\text{pair}}^{2}$, $\sigma_{\text{ldp}}^{2}$, $\sigma_{\text{cdp}}^{2}$ versus (a) privacy budget $\varepsilon$ at $p{=}0.5$ and (b) connectivity $p$ at $\varepsilon{=}10$.
The performance of differentially private D-SGD hinges on the independent portion of each agent’s noise, i.e., the variance that remains after any correlation-based cancellation.
To evaluate our correlated-noise design, we compare $\sigma_{\text{mix}}^{2}$ (the effective independent variance in Whisper D-SGD) against the following state-of-the-art baselines: (i) pairwise correlation, by replacing $\mathbf{R}_{\text{mix}}$ in (10) with $\mathbf{R}_{\text{pair}}=\sigma_{\text{pair}}^{2}\mathbf{I}_{n}+\sigma_{\text{cor}}^{2}\mathbf{L}$, (ii) LDP, with $\mathbf{R}_{\text{ldp}}=\sigma_{\text{ldp}}^{2}\mathbf{I}_{n}$, and (iii) CDP: $\sigma_{\text{cdp}}^{2}=\sigma_{\text{ldp}}^{2}/{n}$.
Figure 1 reports solutions of Problem (10), comparing $\sigma_{\text{mix}}^{2}$, $\sigma_{\text{pair}}^{2}$, $\sigma_{\text{ldp}}^{2}$, and $\sigma_{\text{cdp}}^{2}$ under varying privacy budgets $\varepsilon$ (Fig. 1(a)) and network connectivity $p$ in an Erdős–Rényi graph (Fig. 1(b)).
The fully independent LDP noise imposes the highest variance $\sigma_{\text{ldp}}^{2}$, as it forgoes any correlation-based cancellation (it serves as a practical upper bound). Conversely, $\sigma_{\text{cdp}}^{2}$ acts as a theoretical lower bound, representing the ideal CDP-level variance. As $\varepsilon$ and $p$ increase (weaker privacy requirements or denser connectivity), correlation-based schemes ($\sigma_{\text{pair}}^{2}$ and $\sigma_{\text{mix}}^{2}$) become more effective, with Whisper D-SGD consistently achieving $\sigma_{\text{mix}}^{2}\ll\sigma_{\text{pair}}^{2}$. This gap is particularly large for sparser networks (lower $p$), since $\mathbf{R}_{\text{mix}}$ leverages the entire mixing matrix $\mathbf{W}$, rather than just the graph Laplacian $\mathbf{L}$. Overall, Whisper D-SGD brings LDP-level privacy (through $\sigma_{\text{mix}}^{2}$) much closer to the ideal CDP regime ($\sigma_{\text{cdp}}^{2}$).
VI Experimental Evaluation
(a) Synthetic, $p=0.5$.
(b) Synthetic, $\varepsilon=10$.
Figure 2: Quadratic Optimization. Optimality gap versus (a) privacy budget $\varepsilon$ at $p=0.5$ and (b) connectivity $p$ at $\varepsilon=10$.
(c) a9a dataset, $p=0.5$.
(d) a9a dataset, $\varepsilon=10$.
Figure 3: Logistic Regression. Test loss versus (a) privacy budget $\varepsilon$ at $p=0.5$ and (b) connectivity $p$ at $\varepsilon=10$.
We evaluate Whisper D-SGD on synthetic quadratic objectives and real-world logistic regression, under varying privacy budgets $\varepsilon$ and network connectivity $p$. Our code is available at https://github.com/arodio/WhisperDSGD.
VI-A Experimental Setup
VI-A1 Network and DP parameters
We simulate $n=20$ agents on an Erdős–Rényi graph with varying connectivity $p\in\{0.2,0.4,0.6,0.8,1.0\}$. We adopt Metropolis–Hastings weights $w_{ij}=\tfrac{1}{|\mathcal{N}_{i}|+1}$ if $j\in\mathcal{N}_{i}$ [19]. Following standard benchmarks [18], we set DP parameters $\varepsilon\in\{3,5,7,10,15,20,25,30,40\}$, $\delta=10^{-5}$, and clipping threshold $C=0.1$.
VI-A2 Algorithms
We compare (i) LDP [13, 14], which adds independent local noise; (ii) DECOR [18], representative of pairwise correlation; and (iii) Whisper D-SGD, our mixing-based correlation approach. All algorithms run for $T=5000$ iterations. Results are averaged over 10 random seeds.
VI-A3 Tasks
We consider two machine learning tasks:
(i) Quadratic Optimization with strictly convex objectives:
$f_{i}(x_{1},x_{2})=\begin{cases}15(x_{1}+i)^{2}+x_{2}^{2},&i=1,\dots,\tfrac{n}{2},\\ 15(x_{1}-i)^{2}+x_{2}^{2},&i=\tfrac{n}{2}+1,\dots,n.\end{cases}$ (11)
To introduce data heterogeneity, we rotate $f_{i}$, $i=\tfrac{n}{2}+1,\dots,n$ around the local minimizers $(i,0)$ by an angle $\theta=15^{\circ}$ [25]. We use a diminishing step size $\eta_{t}=1/\sqrt{t}$, with $\eta_{1}=10^{-2}$.
(ii) Logistic Regression. We train a regularized linear model on the a9a LIBSVM dataset [26], with an 80-20 train-test split. We simulate data heterogeneity by randomly partitioning samples among agents using a Dirichlet distribution with parameter 10 [18]. We tune the step-size in $\{10^{-1},5\cdot 10^{-2},10^{-2},5\cdot 10^{-3},10^{-3}\}$ and fix batch size to 128.
VI-B Experimental Results
Figures 33 show the empirical privacy-utility trade-off on both (i) the quadratic optimization task (reporting optimality gap $\frac{1}{n}\!\sum_{i=1}^{n}f_{i}(x_{i}^{(T)})-f^{*}$) and (ii) the real-world logistic regression task (reporting test loss). Panel (a) of each figure varies the privacy budget $\varepsilon$ under fixed connectivity $p=0.5$, while panel (b) varies $p$ under fixed $\varepsilon=10$.
Across all $\varepsilon$ values, both correlation-based approaches (DECOR and Whisper D-SGD) significantly outperform LDP, decreasing final losses by up to an order of magnitude (Figs. 2(a)2(c)). This confirms that correlated noise effectively mitigates the performance drop caused by local noise.
Whisper D-SGD consistently outperforms DECOR for all network connectivities (Figs. 2(b)2(d)). For fully connected networks ($p=1$), $\mathbf{W}$ and $\mathbf{L}$ become structurally equivalent, therefore both methods achieve nearly the same utility. However, at lower connectivities ($p\leq 0.4$), Whisper D-SGD outperforms DECOR by the largest margin, confirming the importance of $\mathbf{W}$ in the correlated-noise design.
In summary, Whisper D-SGD significantly enhances the utility–privacy trade-off, particularly in sparse or moderately connected networks. Our results demonstrate the benefits of mixing-aware correlation compared to baseline approaches that disregard network connectivity and mixing weights.
VII Conclusion
In decentralized learning systems, sensitive information about training data can be exposed through shared local models. While LDP mitigates such privacy risks, it suffers from reduced utility compared to CDP due to accumulated local privacy noise. To address this challenge, we propose the Whisper D-SGD algorithm, which generates correlated noise across agents using a covariance-based approach. This method incorporates the mixing matrix into the design of correlated noise, providing strong privacy guarantees with minimal impact on utility. Both theoretical analysis and empirical results demonstrate that Whisper D-SGD improves the privacy–utility trade-off compared to existing methods (e.g., LDP, pairwise correlation) under various privacy budgets, particularly in sparse network topologies. Future directions include extending this framework to directed graphs and jointly optimizing mixing weights and the covariance matrix to improve convergence and further enhance noise cancellation.