### v2

# Optimizing Privacy-Utility Trade-off in Decentralized Learning with Generalized Correlated Noise

Abstract
Decentralized learning enables distributed agents to collaboratively train a shared machine learning model without a central server, through local computation and peer-to-peer communication. Although each agent retains its dataset locally, sharing local models can still expose private information about the local training datasets to adversaries. To mitigate privacy attacks, a common strategy is to inject random artificial noise at each agent before exchanging local models between neighbors. However, this often leads to utility degradation due to the negative effects of cumulated artificial noise on the learning algorithm. In this work, we introduce CorN-DSGD, a novel covariance-based framework for generating correlated privacy noise across agents, which unifies several state-of-the-art methods as special cases. By leveraging network topology and mixing weights, CorN-DSGD optimizes the noise covariance to achieve network-wide noise cancellation. Experimental results show that CorN-DSGD cancels more noise than existing pairwise correlation schemes, improving model performance under formal privacy guarantees.
I Introduction
Training machine learning models traditionally involves centralizing datasets on a single server, which raises scalability and privacy risks [1]. To address these issues, federated learning allows distributed agents to retain their data locally while sharing only data-dependent computations (e.g., local gradients or models) with a central server [2, 3]. Decentralized learning further removes the need for a central server by allowing agents to update their local models and mix them directly with neighbors according to predefined mixing weights [4, 5, 6, 7]. Although federated and decentralized learning algorithms avoid sharing raw data over the network, local models may still expose sensitive information to adversaries through membership inference or gradient inversion attacks [8].
A widely adopted theoretical concept for mitigating such threats is differential privacy (DP), which provides formal privacy guarantees by injecting random noise into data-dependent computations [9]. While the privacy-utility trade-off has been extensively studied in the central DP (CDP) setting–where a server injects noise on the central model [10, 11, 12]–adapting DP to decentralized settings proves substantially more challenging. Under the local DP (LDP) setting [13, 14], each agent individually adds noise to its local model before mixing, protecting its local dataset from adversaries. However, since each agent adds noise independently without coordination, the utility of LDP degrades in terms of model performance [15].
Recent works on differentially private decentralized learning aim to improve LDP utility by combining (i) a small amount of independent noise per agent, sufficient to ensure local privacy, and (ii) a large amount of correlated noise across agents, designed to cancel during mixing [16, 17, 18]. In principle, this cancellation design preserves local privacy while reducing the overall noise after mixing, thereby improving the accuracy of the learned model. Among these works, [16, 17] rely on zero-sum correlation, which is well-suited for addressing honest-but-curious internal neighbors, but ineffective against external eavesdroppers. Meanwhile, [18] proposes pairwise-canceling correlated noise, which restricts noise correlation to pairs of neighboring agents, overlooking how local models are mixed across the network and limiting the potential for broader correlation design.
In this paper, we argue that extending beyond pairwise, neighbor-only correlations can significantly improve the privacy-utility trade-off in decentralized learning. The key contributions are summarized as follows:
  • We present a novel, covariance-based framework for the design of correlated noise across agents, that recovers state-of-the-art approaches as special cases;
  • We generalize existing analyses on the privacy-utility trade-off, showing that the optimal covariance matrix should account for the network topology and mixing weights. Building on this insight, we propose the CorN-DSGD algorithm, which optimizes the noise covariance to achieve network-wide noise cancellation.
Experiments with various privacy budgets and network connectivity levels show that CorN-DSGD achieves superior privacy-utility trade-offs compared to pairwise approaches, especially in weakly connected networks.
II Privacy-Preserving Decentralized Learning
We consider $n$ agents, $\mathcal{V}=\{1,\dots,n\}$, each with a local dataset $D_{i}$, aiming to learn the parameters $x\in\mathbb{R}^{d}$ of a shared machine learning model, by minimizing the global objective:
$F(x)\triangleq\frac{1}{n}\sum_{i=1}^{n}\left[F_{i}(x)\triangleq\frac{1}{|D_{i}|}\sum_{\xi_{i}\in D_{i}}\ell(x,\xi_{i})\right],$ (1)
where $\ell(x,\xi_{i})$ is the loss of parameter $x$ on sample $\xi_{i}\in D_{i}$ and $F_{i}(x)$ is the local objective, known only to agent $i$. Agents communicate over a network modeled by an undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where an edge $(i,j)\in\mathcal{E}$ indicates that agents $i$ and $j$ are neighbors, i.e., they can directly communicate. For simplicity, we present the case of scalar parameters ($d=1$), with straightforward extension to vector parameters.
II-A Decentralized Stochastic Gradient Descent (DSGD)
Problem (1) is commonly solved via decentralized optimization algorithms like Decentralized Stochastic Gradient Descent (DSGD) [6]. Each iteration $t\in\{1,\dots,T\}$ involves two steps:
a) Local step: Each agent $i$ computes a stochastic gradient $g_{i}^{(t)}=\ell^{\prime}(x_{i}^{(t)},\xi_{i}^{(t)})$, where $\xi_{i}^{(t)}$ is a data point sampled from agent $i$’s dataset $D_{i}$, and then updates its parameter:
$\textstyle x_{i}^{(t+\frac{1}{2})}=x_{i}^{(t)}-\eta_{t}g_{i}^{(t)},$ (2a)
where $\eta_{t}$ is the step-size.
b) Mixing step: Agents exchange parameters with neighbors over the network and compute a weighted average:
$\textstyle{x}_{i}^{(t+1)}=\sum_{j=1}^{n}w_{ij}x_{j}^{(t+\frac{1}{2})},$ (2b)
where $w_{ij}\triangleq[\mathbf{W}]_{ij}$ are mixing weights defined by the mixing matrix $\mathbf{W}\in\mathbb{R}^{n\times n}$, with $w_{ij}=0$ whenever $(i,j)\notin\mathcal{E}$. Convergence conditions for DSGD require $\mathbf{W}$ to be doubly stochastic ($\mathbf{W}\mathbf{1}=\mathbf{1}$ and $\mathbf{1}^{\top}\mathbf{W}=\mathbf{1}^{\top}$) and the second largest eigenvalue $\lambda_{2}(\mathbf{W}^{\top}\mathbf{W})$ strictly below one [19, Section II.B].
II-B Differentially-Private DSGD
In decentralized learning, we aim to protect agents’ local models from potential adversaries, which can be either external eavesdroppers or curious internal agents. For clarity of exposition, we address external eavesdroppers in the remainder of the paper and defer the discussion on honest-but-curious agents to Section V-D. Our privacy goal is to prevent the adversary from inferring whether any specific agent—and by extension, its entire local dataset—is participating in the training process. [^1] An alternative, weaker privacy objective is to protect each individual data sample rather than the entire dataset of an agent. Our framework extends to this sample-level DP setting in a straightforward manner.
Two sets of datasets $D=\{D_{1},D_{2},\dots,D_{n}\}$ and $D^{\prime}=\{D^{\prime}_{1},D^{\prime}_{2},\dots,D^{\prime}_{n}\}$ are agent-level neighbors if they differ by exactly one agent’s data ($D_{j}\neq D^{\prime}_{j}$ for some $j$, and $D_{i}=D^{\prime}_{i}$ for all $i\neq j$). A randomized mechanism $M(\cdot)$ is agent-level $(\varepsilon,\delta)$-DP if, for all neighboring datasets $D,D^{\prime}$ and all subsets $S$ of outputs, we have:
$\mathrm{Pr}\bigl{[}M(D)\in S\bigr{]}\leq e^{\varepsilon}\mathrm{Pr}\bigl{[}M(D^{\prime})\in S\bigr{]}+\delta.$ (3)
In the absence of a central server, privacy guarantees must be enforced locally by each agent. A classical LDP mechanism is presented in [13]. At each iteration $t$, each agent clips its own local gradient: $\hat{g}_{i}^{(t)}=\min\{1,\frac{C}{|g_{i}^{(t)}|}\}\cdot g_{i}^{(t)}$. It then adds noise locally and independently: $\tilde{g}_{i}^{(t)}=\hat{g}_{i}^{(t)}+v_{i}^{(t)}$. Finally, it applies the mixing step: $\tilde{x}_{i}^{(t+1)}=\sum_{j=1}^{n}w_{ij}(\tilde{x}_{j}^{(t)}-\eta_{t}\tilde{g}_{j}^{(t)})$. Specifically, the LDP Gaussian mechanism adds noise $v_{i}^{(t)}\sim\mathcal{N}(0,\sigma_{\text{ldp}}^{2})$, such that, for sufficiently large noise variance $\sigma_{\text{ldp}}^{2}$, all local models $\{\tilde{x}_{i}^{(t+1)}\}_{i=1}^{n}$ are agent-level $(\varepsilon,\delta)$-LDP [14].
Despite its strong privacy guarantees, this LDP mechanism suffers from utility degradation, since the added independent privacy noise accumulates over the network, leading to higher variance in the mixed models, slower convergence, and reduced model performance [15]. Recognizing the utility degradation inherent to LDP, prior work [18] proposed pairwise-canceling correlated noise across agents. In DECOR [18], each agent adds noise $v_{i}^{(t)}=u_{i}^{(t)}+\sum_{j\in\mathcal{N}_{i}}v_{ij}^{(t)}$, where $u_{i}^{(t)}\sim\mathcal{N}(0,\sigma_{\text{pair}}^{2})$ is independent local noise, and each pair of neighbors $(i,j)\in\mathcal{E}$ shares a correlated noise term $v_{ij}^{(t)}=-v_{ji}^{(t)}\sim\mathcal{N}(0,\sigma_{\text{cor}}^{2})$ that cancels in the mixing step. Distinct pairs of neighbors produce mutually uncorrelated components, so that $v_{ij}^{(t)}$ is independent of $v_{k\ell}^{(t)}$ whenever $\{i,j\}\cap\{k,\ell\}=\emptyset$. However, DECOR restricts noise correlations to neighboring agents and offers fewer degrees of freedom in noise design, particularly in sparse networks.
This paper introduces a novel covariance-based framework for generating correlated noise across agents, showing how the correlation structure should be designed by taking into account the effects of network topology and mixing weights.
III A Covariance-Based Framework for Generating Correlated Noise Across Agents
We reinterpret the differentially private DSGD framework from a noise-covariance perspective. Instead of drawing noise components $\{v_{i}^{(t)}\}_{i=1}^{n}$ independently or from pairwise correlations, we allow agents to sample from a multivariate Gaussian:
$\mathbf{v}^{(t)}\;=\;\bigl{[}v_{1}^{(t)},\dots,v_{n}^{(t)}\bigr{]}^{\top}\;\sim\;\mathcal{N}\!\bigl{(}\mathbf{0},\mathbf{R}\bigr{)},$ (4)
where $\mathbf{R}\in\mathbb{R}^{n\times n}$ is an arbitrary covariance matrix. By defining $\tilde{\mathbf{x}}^{(t)}\triangleq[\tilde{x}_{1}^{(t)},\dots,\tilde{x}_{n}^{(t)}]^{\top}$ and $\tilde{\mathbf{g}}^{(t)}=[\tilde{g}_{1}^{(t)},\ldots,\tilde{g}_{n}^{(t)}]^{\top}$, the perturbed model update at iteration $t$ becomes:
$\tilde{\mathbf{x}}^{(t+1)}=\mathbf{W}\left(\tilde{\mathbf{x}}^{(t)}-\eta_{t}\tilde{\mathbf{g}}^{(t)}\right).$ (5)
In practice, each agent can generate $\mathbf{v}^{(t)}$ locally by sharing $\mathbf{R}$ and a random seed $s$, drawing $\mathbf{s}^{(t)}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{n})$ independently using the seed $s$, and locally computing $\mathbf{v}^{(t)}=\mathbf{R}^{1/2}\mathbf{s}^{(t)}$.
This covariance-based approach offers several advantages: (i) it allows for more flexible correlation structures with noise cancellation beyond immediate neighbors; (ii) it is directly applicable to directed graphs, unlike pairwise correlation; and (iii) it recovers state-of-the-art approaches as special cases:
  1. 1.
    LDP. With independent noise, the covariance matrix is $\mathbf{R}=\sigma_{\text{ldp}}^{2}\mathbf{I}_{n}$, where $\mathbf{I}_{n}\in\mathbb{R}^{n\times n}$ is the identity matrix;
  2. 2.
    DECOR. The independent noise term $u_{i}^{(t)}\sim\mathcal{N}(0,\sigma_{\text{pair}}^{2})$ contributes $\sigma_{\text{pair}}^{2}\mathbf{I}_{n}$ to $\mathbf{R}$. In addition, each edge $(i,j)\in\mathcal{E}$ adds a pairwise correlated component $v_{ij}^{(t)}=-v_{ji}^{(t)}$ with variance $\sigma_{\text{cor}}^{2}$. In matrix form, each edge contributes $\sigma_{\text{cor}}^{2}(\mathbf{e}_{i}-\mathbf{e}_{j})(\mathbf{e}_{i}-\mathbf{e}_{j})^{\top}$, where $\mathbf{e}_{i}$ is the $i$-th standard basis vector, and summing over all edges yields $\sigma_{\text{cor}}^{2}\mathbf{L}$, where $\mathbf{L}=\sum_{(i,j)\in\mathcal{E}}(\mathbf{e}_{i}-\mathbf{e}_{j})(\mathbf{e}_{i}-\mathbf{e}_{j})^{\top}$ is the undirected graph Laplacian. The total covariance is $\mathbf{R}=\sigma_{\text{pair}}^{2}\mathbf{I}_{n}+\sigma_{\text{cor}}^{2}\mathbf{L}$.
Constraining $\mathbf{R}$ to the graph Laplacian $\sigma_{\text{cor}}^{2}\mathbf{L}$ limits noise correlation, particularly in sparse graphs. For instance, in a star topology, $\mathbf{R}=\sigma_{\text{pair}}^{2}\mathbf{I}_{n}+\sigma_{\text{cor}}^{2}\mathbf{L}$ forces $n(n-1)-2(n-1)$ covariance entries to zero, and overlooks all leaf-to-leaf noise cancellations, which a general $\mathbf{R}\succeq 0$ can instead exploit.
IV Privacy-Utility Analysis
The covariance matrix $\mathbf{R}$ affects both utility (i.e., convergence rate) and privacy (i.e., $(\varepsilon,\delta)$-LDP guarantees).
IV-A Utility Analysis
We introduce a virtual sequence $\hat{\mathbf{x}}^{(t+1)}=\mathbf{W}(\tilde{\mathbf{x}}^{(t)}-\eta_{t}\hat{\mathbf{g}}^{(t)})$, which begins at the noisy iterate $\tilde{\mathbf{x}}^{(t)}$ but applies a noise-free DSGD step. This construction isolates the error introduced by the privacy noise and simplifies the analysis in [18, Thm. 2].
Assumption 1.
At iteration $t$, the privacy noise $\mathbf{v}^{(t)}$ is independent of the noise-free state $\hat{\mathbf{x}}^{(t+1)}$ and prior states $\mathcal{F}^{(t)}=\{(\mathbf{v}^{(1)},\hat{\mathbf{x}}^{(2)}),\dots,(\mathbf{v}^{(t-1)},\hat{\mathbf{x}}^{(t)})\}$: $\mathbb{E}[\mathbf{v}^{(t)}|{\hat{\mathbf{x}}^{(t+1)}},\mathcal{F}^{(t)}]=\mathbf{0}.$
Theorem 1.
Under Assumption 1, the expected squared error between consecutive iterates of our covariance-based framework satisfies:
$\mathbb{E}\Bigl{[}\bigl{\|}\tilde{\mathbf{x}}^{(t+1)}-\tilde{\mathbf{x}}^{(t)}\bigr{\|}^{2}\;\Big{|}\;\mathcal{F}^{(t)}\Bigr{]}$
$=\;\underbrace{\mathbb{E}\Bigl{[}\bigl{\|}\hat{\mathbf{x}}^{(t+1)}-\tilde{\mathbf{x}}^{(t)}\bigr{\|}^{2}\;\Big{|}\;\mathcal{F}^{(t)}\Bigr{]}}_{\begin{subarray}{c}\text{Noise-free DSGD error}\\ \text{with clipped gradients}\end{subarray}}\;+\;\underbrace{\eta_{t}^{2}\,\mathrm{Tr}\!\bigl{(}\mathbf{W}\mathbf{R}\mathbf{W}^{\top}\bigr{)}}_{\text{Privacy-noise variance}}.$ (6)
Proof Sketch.
First, decompose $\|\tilde{\mathbf{x}}^{(t+1)}-\tilde{\mathbf{x}}^{(t)}\|^{2}$ into the sum of $\|\hat{\mathbf{x}}^{(t+1)}-\tilde{\mathbf{x}}^{(t)}\|^{2}$ and $\|\tilde{\mathbf{x}}^{(t+1)}-\hat{\mathbf{x}}^{(t+1)}\|^{2}$. Taking conditional expectation, the cross-term vanishes by Assumption 1. Next, observe $\|\tilde{\mathbf{x}}^{(t+1)}-\hat{\mathbf{x}}^{(t+1)}\|^{2}=\eta_{t}^{2}\|\mathbf{W}\mathbf{v}^{(t)}\|^{2}$. Finally, $\mathbf{R}=\mathbb{E}[\mathbf{v}^{(t)}(\mathbf{v}^{(t)})^{\top}]$ yields the variance term $\eta_{t}^{2}\,\mathrm{Tr}(\mathbf{W}\mathbf{R}\mathbf{W}^{\top})$. ∎
Theorem 1 shows that each iteration of our covariance-based framework incurs a baseline DSGD error (independent of $\mathbf{R}$) plus a variance term proportional to $\mathrm{Tr}(\mathbf{W}\mathbf{R}\mathbf{W}^{\top})$.
Our analysis captures the interplay between $\mathbf{W}$ and $\mathbf{R}$, explaining how correlated noise propagates and potentially cancels across agents. Prior analyses overlooked this mutual dependency, which is central to our covariance-based view.
Additionally, the variance $\mathrm{Tr}(\mathbf{W}\mathbf{R}\mathbf{W}^{\top})$ is linear in $\mathbf{R}$: smaller $[\mathbf{R}]_{ii}$ components speed-up convergence. This implies that the optimal correlation structure should depend on the underlying network topology and mixing weights.
IV-B Privacy Analysis
We provide $(\varepsilon,\delta)$-LDP guarantees for our framework with arbitrary $\mathbf{R}\succ 0$. Theorem 2 bounds $\varepsilon$ in terms of number of iterations $T$, clipping threshold $C$, and inverse covariance $\mathbf{R}^{-1}$.
Theorem 2.
Our covariance-based framework is ($\varepsilon,\delta$)-LDP:
$\textstyle\varepsilon\leq 2C^{2}T\max\limits_{i\in\mathcal{V}}[\mathbf{R}^{-1}]_{ii}+2C\sqrt{2T\log(\frac{1}{\delta})\max\limits_{i\in\mathcal{V}}[\mathbf{R}^{-1}]_{ii}}.$ (7)
Proof Sketch.
We adapt existing Rényi DP (RDP) arguments from [12, 18] to arbitrary $\mathbf{R}\succ 0$. At each iteration, our framework satisfies $(\alpha,\alpha\varepsilon_{\text{step}})$-RDP with $\varepsilon_{\text{step}}=2C^{2}\max_{i\in\mathcal{V}}[\mathbf{R}^{-1}]_{ii}$ [18, Thm. 6]. Since RDP composes additively over $T$ iterations [12, Prop. 1], the final iterate $\tilde{\mathbf{x}}^{(T)}$ satisfies ($\alpha,T\alpha\varepsilon_{\text{step}}$)-RDP. Converting RDP to $(\varepsilon,\delta)$-DP introduces an additional term $\ln(1/\delta)/(\alpha-1)$ [12, Prop. 3]. Optimizing the expression over $\alpha>1$ concludes the proof. ∎
Theorem 2 directly relates the diagonal elements of $\mathbf{R}^{-1}$, specifically $\max_{i\in\mathcal{V}}[\mathbf{R}^{-1}]_{ii}$, to the privacy budget $\varepsilon$. This result immediately recovers well-known LDP bounds (e.g., [12, Prop. 7]) when $\mathbf{R}=\sigma_{\text{ldp}}^{2}\mathbf{I}_{n}$: larger $[\mathbf{R}]_{ii}$ (i.e., larger variance of privacy noise) implies smaller $[\mathbf{R}^{-1}]_{ii}$, thus smaller $\varepsilon$ (that is, strengthened $(\varepsilon,\delta)$-LDP guarantees).
Invertibility ($\mathbf{R}\succ 0$) is a necessary condition for LDP guarantees under correlated noise: if $\mathbf{R}$ were singular, linear combinations of noise could cancel completely across agents, directly exposing their local models after the mixing step.
V The “CorN-DSGD” Algorithm
(a) $n=20$, $p=0.5$.
(b) $n=20$, $\varepsilon=10$.
(c) $p=0.5$, $\varepsilon=10$.
Figure 1: $\mathrm{Tr}(\mathbf{W}\mathbf{R}_{\text{mix}}\mathbf{W}^{\top})$, $\mathrm{Tr}(\mathbf{W}\mathbf{R}_{\text{pair}}\mathbf{W}^{\top})$, $\mathrm{Tr}(\mathbf{W}\mathbf{R}_{\text{ldp}}\mathbf{W}^{\top})$ vs. varying: (a) privacy budget $\varepsilon$ at $p=0.5$, (b) connectivity $p$ at $\varepsilon=10$, and (c) number of agents $n$ at $p=0.5$, $\varepsilon=10$.
(d) $n=20$, $p=0.5$, $\varepsilon=10$.
Figure 2: $\mathrm{Tr}(\mathbf{W}\mathbf{R}_{\text{mix}}\mathbf{W}^{\top})$ vs. number of independent seeds.
V-A Privacy-Utility Trade-Off
Theorems 1 and 2 together highlight a fundamental tension between utility and privacy: smaller diagonal entries $[\mathbf{R}]_{ii}$ reduce the model-update variance $\mathrm{Tr}(\mathbf{W}\mathbf{R}\mathbf{W}^{\top})$ in Eq. (6) (expediting convergence), yet they increase $[\mathbf{R}^{-1}]_{ii}$ and thus amplify $\varepsilon$ in Eq. (7) (weaker privacy). Conversely, making $[\mathbf{R}]_{ii}$ large improves privacy but slows convergence. A careful choice of $\mathbf{R}$ is therefore crucial to optimize this trade-off. From our analysis, we draw the following guidelines:
(G1) Invertibility. A necessary condition for $(\varepsilon,\delta)$-LDP is $\mathbf{R}\succ 0$ (Theorem 2). For this reason, we partition $\mathbf{R}$ into an independent noise component ($\sigma_{\text{mix}}^{2}\mathbf{I}_{n}$), that ensures invertibility, and an arbitrary correlated component ($\mathbf{R}_{\text{cor}}$):
$\mathbf{R}_{\text{mix}}=\sigma_{\text{mix}}^{2}\mathbf{I}_{n}+\mathbf{R}_{\text{cor}}.$ (8)
The covariance $\mathbf{R}_{\text{mix}}$ in (8) generalizes prior approaches as special cases: (i) LDP: $\mathbf{R}_{\text{cor}}=\mathbf{0}$; (ii) DECOR: $\mathbf{R}_{\text{cor}}=\sigma_{\text{cor}}^{2}\mathbf{L}$. We remark $\sigma_{\text{mix}}^{2}>0$; if not, $\mathbf{R}_{\text{mix}}$ would be singular, and $\mathbf{v}\sim\mathcal{N}(\mathbf{0},\mathbf{R}_{\text{mix}})$ could make $\mathbf{W}\mathbf{v}=\mathbf{0}$, i.e., no privacy noise.
(G2) Optimization Problem. We optimize $\mathbf{R}_{\text{mix}}$ to balance utility (Theorem 1) and $(\varepsilon,\delta)$-LDP privacy (Theorem 2):
$\underset{\mathbf{R}_{\text{mix}}\succ 0}{\text{minimize}}\quad$ $\mathrm{Tr}\bigl{(}\mathbf{W}\mathbf{R}_{\text{mix}}\mathbf{W}^{\top}\bigr{)}$ (9a)
subject to $[\mathbf{R}_{\text{mix}}^{-1}]_{ii}\leq\frac{\varepsilon^{2}}{16C^{2}T\log(1/\delta)},~{}\forall\,i\in\mathcal{V}.$ (9b)
Problem (9) can be formulated as a convex semidefinite program (SDP). The objective is linear in $\mathbf{R}_{\text{mix}}$, and the mapping $\mathbf{R}_{\text{mix}}\mapsto\mathbf{R}_{\text{mix}}^{-1}$ is matrix-convex on $\mathbf{R}_{\text{mix}}\succ 0$ [20, §4.6.2]. By applying Schur-complement constraints, each bound on $[\mathbf{R}_{\text{mix}}^{-1}]_{ii}$ becomes a linear matrix inequality. Thus, Problem (9) can be solved efficiently by standard SDP solvers (e.g., MOSEK [21]). Also, our SDP formulation is scalable, as we verify numerically with networks of up to 100 agents (Fig. 1(c)). Solving (9) yields the optimal solutions $\sigma_{\text{mix}}^{2,\star}$ and $\mathbf{R}_{\text{cor}}^{\star}$, from which we compute $\mathbf{R}_{\text{mix}}^{\star}=\sigma_{\text{mix}}^{2,\star}\mathbf{I}_{n}+\mathbf{R}_{\text{cor}}^{\star}$.
Proposition 1.
Problem (9) attains the largest utility among all covariances satisfying (8); every such solution is $(\varepsilon,\delta)$-LDP.
Proof.
First, the objective in (9a) minimizes the privacy-noise variance in Theorem 1, Eq. (6). Second, applying Eq. (9b) to the privacy bound in Theorem 2, Eq. (7) shows that our framework is ($\varepsilon^{*},\delta$)-LDP, where $\varepsilon^{*}\leq\frac{\varepsilon^{2}}{8\log(1/\delta)}+\frac{\varepsilon}{\sqrt{2}}\leq\varepsilon$. ∎
V-B Correlated Noise DSGD (CorN-DSGD)
Algorithm 1 CorN-DSGD
1:Input: mixing matrix $\mathbf{W}$, $(\varepsilon,\delta)$-DP parameters, number of iterations $T$, clipping threshold $C$, random seed $s$.
2:Initialize $\!\mathbf{R}_{\text{mix}}=\sigma_{\text{mix}}^{2}\mathbf{I}_{n}+\mathbf{R}_{\text{cor}}$triangleright guideline (G1)
3:Optimize (9) to find $\mathbf{R}_{\text{mix}}^{\star}$triangleright guideline (G2)
4:Share ($\mathbf{R}_{\text{mix}}^{\star}$, $s$) among all agents
5:for $t$ in $10\dots T-$ 1 do
6:     $\hat{\mathbf{g}}^{(t)}=\mathrm{clip}\bigl{(}\mathbf{g}^{(t)};C\bigr{)}$triangleright clipped gradient
7:     $\mathbf{v}^{(t)}=\mathbf{R}_{\text{mix}}^{\star 1/2}\mathbf{s}^{(t)}$, $\mathbf{s}^{(t)}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{n})$triangleright privacy noise
8:     $\tilde{\mathbf{g}}^{(t)}=\hat{\mathbf{g}}^{(t)}+\mathbf{v}^{(t)}$triangleright perturbed gradient
9:     $\tilde{\mathbf{x}}^{(t+1)}=\mathbf{W}\left(\tilde{\mathbf{x}}^{(t)}-\eta_{t}\tilde{\mathbf{g}}^{(t)}\right)$triangleright perturbed update
10:end for
11:return $\tilde{\mathbf{x}}^{(T)}$
We now present CorN-DSGD (Algorithm 1), a novel differentially private DSGD algorithm that integrates our covariance-based framework with the theoretical guidelines.
CorN-DSGD: (i) initializes $(\sigma_{\text{mix}}^{2},\mathbf{R}_{\text{cor}})$ so that $\mathbf{R}_{\text{mix}}$ is invertible (line 2); (ii) optimizes Problem (9) via standard SDP solvers (line 3); (iii) shares the optimal $\mathbf{R}_{\text{mix}}^{\star}$ and a global random seed $s$ among all agents (line 4); (iv) executes the standard differentially-private DSGD algorithm (lines 510), except agents add privacy noise $\mathbf{v}^{(t)}=\mathbf{R}_{\text{mix}}^{\star 1/2}\mathbf{s}^{(t)}$, where $\mathbf{s}^{(t)}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{n})$ is locally generated from seed $s$ (line 7).
Once agents share $\mathbf{R}_{\text{mix}}^{\star}$ and seed $s$, CorN-DSGD runs fully decentralized, incurring no extra communication overhead than baseline DSGD (lines 510). However, computing $\mathbf{R}_{\text{mix}}^{\star}$ requires either a centralized or a distributed procedure (lines 24). In the centralized approach, a single node with knowledge of $\mathbf{W}$ and parameters $(\varepsilon,\delta,T,C)$ solves (9) and broadcasts $(\mathbf{R}_{\text{mix}}^{\star},s)$ to all agents. Alternatively, a distributed setup involves an initial consensus phase on $\mathbf{W}$ and $s$, then each agent locally solves (9). This approach removes reliance on any single node but incurs an initial communication overhead.
V-C Comparison with Related Works
(a) Synthetic, $p=0.5$.
(b) Synthetic, $\varepsilon=10$.
Figure 3: Quadratic Optimization. Optimality gap versus (a) privacy budget $\varepsilon$ at $p=0.5$ and (b) connectivity $p$ at $\varepsilon=10$.
(c) a9a dataset, $p=0.5$.
(d) a9a dataset, $\varepsilon=10$.
Figure 4: Logistic Regression. Test loss versus (a) privacy budget $\varepsilon$ at $p=0.5$ and (b) connectivity $p$ at $\varepsilon=10$.
(a) MNIST, $p=0.5$.
(b) MNIST, $\varepsilon=10$.
Figure 5: Image Classification. Test accuracy versus (a) privacy budget $\varepsilon$ at $p=0.5$ and (b) connectivity $p$ at $\varepsilon=10$.
We compare $\mathrm{Tr}(\mathbf{W}\mathbf{R}_{\text{mix}}\mathbf{W}^{\top})$ (the effective noise variance in CorN-DSGD) against the following state-of-the-art baselines: (i) pairwise correlation, by replacing $\mathbf{R}_{\text{mix}}$ in (9) with $\mathbf{R}_{\text{pair}}=\sigma_{\text{pair}}^{2}\mathbf{I}_{n}+\sigma_{\text{cor}}^{2}\mathbf{L}$, and (ii) LDP, with $\mathbf{R}_{\text{ldp}}=\sigma_{\text{ldp}}^{2}\mathbf{I}_{n}$.
Figure 2 reports solutions of Problem (9) for $\mathbf{R}_{\text{mix}}$, $\mathbf{R}_{\text{pair}}$, and $\mathbf{R}_{\text{ldp}}$ in an Erdős–Rényi graph under varying privacy budgets $\varepsilon$ (Fig. 1(a)), network connectivity $p$ (Fig. 1(b)), and number of agents $n$ (Fig. 1(c)). The fully independent LDP noise imposes the highest variance $\mathrm{Tr}(\mathbf{W}\mathbf{R}_{\text{ldp}}\mathbf{W}^{\top})$, as it forgoes any correlation-based cancellation. As $\varepsilon$ and $p$ increase (weaker privacy requirements or denser connectivity), correlation-based schemes ($\mathbf{R}_{\text{pair}}$ and $\mathbf{R}_{\text{mix}}$) become more effective, and $\mathbf{R}_{\text{mix}}$ consistently achieves lower variance than $\mathbf{R}_{\text{pair}}$. This gap is particularly large for sparser networks (lower $p$), since $\mathbf{R}_{\text{mix}}$ leverages broader covariance structures (through $\mathbf{R}_{\text{cor}}$), rather than just the graph Laplacian $\mathbf{L}$.
V-D Honest-But-Curious (HBC) Setting
We extend Problem (9) to account for honest-but-curious (HBC) agents. Let $\mathcal{S}=\{s_{k}\}_{k=1}^{m}$ be $m$ independent seeds. Seed $s_{k}$ is known to the (possibly overlapping) set of agents $\mathcal{H}_{k}\subseteq\mathcal{V}$. The covariance is decomposed as: $\mathbf{R}_{\text{cor}}=\sum_{k=1}^{m}\mathbf{R}_{\text{cor}}^{(k)}$, with $\mathrm{supp}(\mathbf{R}_{\text{cor}}^{(k)})\subseteq\mathcal{H}_{k}\times\mathcal{H}_{k}$. Thus the pair $(\mathbf{R}_{\text{cor}}^{(k)},s_{k})$ is hidden from every agent outside $\mathcal{H}_{k}$.
For any adversarial coalition $I\subseteq\mathcal{V}$, define the set of unknown seeds $\mathcal{U}(I)\coloneqq\{k:\mathcal{H}_{k}\cap I=\emptyset\}$. The covariance unknown to $I$ is $\mathbf{R}_{\text{mix}}^{(I)}\coloneqq\sigma_{\text{mix}}^{2}\mathbf{I}_{n}+\sum_{k\in\mathcal{U}(I)}\mathbf{R}_{\text{cor}}^{(k)}$. We protect against coalitions of size at most $q$: $\mathcal{C}_{q}\coloneqq\{I\subseteq\mathcal{V}:|I|\leq q\}$.
In this context, the utility analysis in Theorem 1 remains unchanged, while Theorem 2 requires only minor adaptation, as each coalition $I\in\mathcal{C}_{q}$ imposes its own privacy constraint:
$\textstyle\underset{\sigma_{\text{mix}}^{2}>0,\{\mathbf{R}_{\text{cor}}^{(k)}\}}{\text{minimize}}\quad$ $\mathrm{Tr}\bigl{(}\mathbf{W}\mathbf{R}_{\text{mix}}\mathbf{W}^{\top}\bigr{)}$ (10a)
subject to $\textstyle\mathbf{R}_{\text{mix}}=\sigma_{\text{mix}}^{2}\mathbf{I}_{n}+\sum_{k=1}^{m}\mathbf{R}_{\text{cor}}^{(k)},$ (10b)
$\textstyle\mathbf{R}_{\text{cor}}^{(k)}\succeq\mathbf{0},~{}\mathrm{supp}(\mathbf{R}_{\text{cor}}^{(k)})\subseteq\mathcal{H}_{k}\times\mathcal{H}_{k},$ (10c)
$\textstyle[(\mathbf{R}_{\text{mix}}^{(I)})^{-1}]_{ii}\leq\frac{\varepsilon^{2}}{16C^{2}T\log(1/\delta)},$ (10d)
$\forall I\in\mathcal{C}_{q},i\not\in I.$
Problem (10) is solvable in polynomial time. Special cases are: (i) external eavesdropper ($q=0$, $m=1$); (ii) HBC without collusion ($q=1$); (iii) DECOR ($q\geq 1$, $m=|\mathcal{E}|$); (iv) LDP ($q=n-1$, $m=n$). Fig. 2 shows the value of $\mathrm{Tr}(\mathbf{W}\mathbf{R}_{\text{mix}}\mathbf{W}^{\top})$ for different numbers of independent seeds ($m\in\{1,\dots,n\}$) and disjoint groups of agents ($\bigcap_{k}\mathcal{H}_{k}=\emptyset$).
VI Experimental Evaluation
VI-A Experimental Setup
VI-A1 Network and DP parameters
We simulate $n=20$ agents on an Erdős–Rényi graph with varying connectivity $p\in\{0.2,0.4,0.6,0.8,1.0\}$. We use Metropolis–Hastings weights: $w_{ij}=\tfrac{1}{|\mathcal{N}_{i}|+1}$ if $j\in\mathcal{N}_{i}$ [19]. For comparison with [18], we focus on external eavesdroppers observing all agents’ communications ($q=0$, $m=1$), and we set DP parameters $\varepsilon\in\{3,5,7,10,15,20,25,30,40\}$, $\delta=10^{-5}$, $C=0.1$.
VI-A2 Algorithms
We compare (i) LDP [13, 14], which adds independent local noise; (ii) DECOR [18], representative of pairwise correlation; and (iii) CorN-DSGD, our covariance-based approach. All algorithms run for $T=5000$ iterations. Results are averaged over 10 random seeds.
VI-A3 Tasks
We consider three machine learning tasks:
(i) Quadratic Optimization with strictly convex objectives:
$f_{i}(x_{1},x_{2})=\begin{cases}15(x_{1}+i)^{2}+x_{2}^{2},&i=1,\dots,\tfrac{n}{2},\\ 15(x_{1}-i)^{2}+x_{2}^{2},&i=\tfrac{n}{2}+1,\dots,n.\end{cases}$ (11)
To introduce data heterogeneity, we rotate $f_{i}$, $i=\tfrac{n}{2}+1,\dots,n$ around the local minimizers $(i,0)$ by an angle $\theta=15^{\circ}$ [22]. We use a diminishing step size $\eta_{t}=1/\sqrt{t}$, with $\eta_{1}=10^{-2}$.
(ii) Logistic Regression. We train a regularized linear model on the a9a LIBSVM dataset [23], with an 80-20 train-test split. We simulate data heterogeneity by randomly partitioning samples among agents using a Dirichlet distribution with parameter 10 [18]. We tune the step-size in $\{10^{-1},5\cdot 10^{-2},10^{-2},5\cdot 10^{-3},10^{-3}\}$ and fix batch size to 128.
(iii) Image Classification. We train a shallow neural network on the MNIST dataset [24], with the same train-test split, data partitioning scheme, and hyperparameter grid as in task (ii).
VI-B Experimental Results
Figures 45 show the empirical privacy-utility trade-off on (i) the quadratic optimization task (reporting optimality gap $\frac{1}{n}\!\sum_{i=1}^{n}f_{i}(x_{i}^{(T)})-f^{*}$), (ii) logistic regression task (reporting test loss), and (iii) image classification task (reporting test accuracy). Panel (a) of each figure varies the privacy budget $\varepsilon$ under fixed connectivity $p=0.5$, while panel (b) varies $p$ under fixed $\varepsilon=10$. Across all $\varepsilon$ values, both correlation-based approaches (DECOR and CorN-DSGD) outperform LDP (Figs. 45(a)). CorN-DSGD consistently outperforms DECOR for all network connectivities (Figs. 45(b)). For fully connected networks ($p=1$), $\mathbf{R}_{\text{cor}}$ and $\mathbf{L}$ become proportional, therefore both methods achieve nearly the same utility. However, at lower connectivities ($p\leq 0.4$), CorN-DSGD outperforms DECOR by the largest margin, confirming the benefits of our covariance-based framework compared to prior approaches in sparse networks.
VII Conclusion
In serverless decentralized learning, sensitive information about training data can be exposed through shared local models. While LDP mechanisms mitigate such privacy threats, they suffer from reduced utility due to independently added local privacy noise. We propose the CorN-DSGD algorithm, which generates correlated noise across agents using a covariance-based approach. This method incorporates the mixing weights into the design of correlated noise, offering strong privacy guarantees with minimal impact on utility. Both theoretical analysis and empirical results demonstrate that CorN-DSGD improves the privacy-utility trade-off compared to existing methods (LDP, DECOR) under various privacy budgets, particularly in sparse network topologies.